OpenStudy (bookworm14):

Can someone please help me? I have worked this problem out at least 7 times, and still can't get the right answer!!! Can someone please help me work it out??? 12x-3x+11 > 4x-(17-9x) A. x< -11/14 B. x>-7 C. x>-14/11 D. x<7 I keep getting a negative 2, which is not correct :(

OpenStudy (chosenmatt):

need help?

OpenStudy (bookworm14):

yes please @chosenmatt

Directrix (directrix):

12x-3x+11 > 4x-(17-9x) 9x + 11 > 4x - 17 + 9x @Bookworm14 Did you get this far the seven times you worked this?

OpenStudy (bookworm14):

yes

OpenStudy (bookworm14):

wait no i lied, how did you get +9 instead of -9?

Directrix (directrix):

9x + 11 > 4x - 17 + 9x 9x + 11 > 13x - 17 9x - 13x > -17 - 11

Directrix (directrix):

- 1(17-9x) = -1(7) + (-1) * (-9x) = -7 + 9x

OpenStudy (bookworm14):

omg my brain was way over thinking this problem can't believe i missed that step, i knew that was where i was going wrong, but i wasn't sure how lol .... i got the answer i htink it is x<7

Directrix (directrix):

Yes, x<7. You remembered to reverse the order of the inequality symbol when you divided by -4. Most people who make an error on this type problem make it at that point.

OpenStudy (bookworm14):

hmmm yes ive done ALOT of these inequalities, so my brain must be overlooking simple steps lol